OpenStudy (unklerhaukus):

Find the integrating factor to make the following equation exact , and then solve; \[\left(3x+\frac6y\right)\mathrm dx+\left(\frac{x^2}y+\frac{3y}x\right)\mathrm dy\]

OpenStudy (unklerhaukus):

\[\newcommand\p \newcommand \p \pat [2] { \tfrac{\partial #1}{\partial #2} } M =x+6y^{-1} \qquad N=x^2y^{-1}+3x^{-1}y \\ M_y =-6y^{-2} \qquad N_x=2xy^{-1}-3x^{-2}y\\ \pat My\neq\pat Nx\implies\text{ not exact}\quad\]

OpenStudy (unklerhaukus):

\[\newcommand\p \newcommand \p \pat [2] { \tfrac{\partial #1}{\partial #2} } \text{let}\quad R =R(x,y)=x^my^n \\ \qquad\qquad R_x =mx^{m-1}y^n \\ \qquad\qquad R_y =nx^my^{n-1} \\ \\ \overline M= R(x,y)M\qquad\qquad\qquad\overline N= R(x,y)N \\ \\ \,\\ \pat {\overline M}y=\pat {\overline N}x\quad\text{ for exactness}\]

OpenStudy (unklerhaukus):

\[\newcommand\p \newcommand \p \pa [2] { \frac{\partial #1}{\partial #2}} \\ \\ \pa{}y\left(R(x,y)(x+6y^{-1})\right) =\pa{}x\left({R(x,y)}(x^2y^{-1}+3x^{-1}y)\right) \\ R_y(x+6y^{-1})-R(6y^{-2}) =R_x(x^2y^{-1}+3x^{-1}y)+R(2xy^{-1}-3x^{-2}y) \\ R_y(x+6y^{-1})-R_x(x^2y^{-1}+3x^{-1}y) =R(2xy^{-1}-3x^{-2}y+6y^{-2}) \\ nx^my^{n-1}(x+6y^{-1})-mx^{m-1}y^n(x^2y^{-1}+3x^{-1}y) =x^my^n(2xy^{-1}-3x^{-2}y+6y^{-2}) \\ nx^my^n(xy^{-1}+6y^{-2})-mx^my^n(xy^{-1}+3x^{-2}y) =x^my^n(2xy^{-1}-3x^{-2}y+6y^{-2}) \\ \\ n(xy^{-1}+6y^{-2})-m(xy^{-1}+3x^{-2}y)=2xy^{-1}-3x^{-2}y+6y^{-2} \\ \]

OpenStudy (unklerhaukus):

equating coefficients i get n-m=2 n=1 m=1 but i think ive made a mistake because n-m=2 should be n+m=2 the correct integrating factor should be R(x,y)=xy

OpenStudy (unklerhaukus):

oh, right i forget the 3 for some reason ....

OpenStudy (unklerhaukus):

\[M=\color{red}3x+6y^{−1}\]

OpenStudy (unklerhaukus):

... \[(3n-m-2)xy^{-1}+6(n-1)y^{-2} =0\\ \,\\ (3n-m-2)=0\qquad n-1=0 \\ \qquad\qquad\qquad\qquad\qquad n=1 \\ 1-m=0 \\ m=1\]

OpenStudy (unklerhaukus):

thats better

OpenStudy (unklerhaukus):

well spotted @UnkleRhaukus, thanks again, i hope you can finish from here..

OpenStudy (unklerhaukus):

\[\begin{align*} \overline M&=xy(3x+6y^{-1})&\overline N &= xy(x^2y^{-1}+3x^{-1}y) \\ &=3x^2y+6x & &=x^3+3y^2 \\ \\ \overline M_y &=3x^2 &\overline N_x &=3x^2 \end{align*}\] exactness

OpenStudy (unklerhaukus):

i always did say there was nothing wrong with talking to your self, ikr.

OpenStudy (unklerhaukus):

Now integrate M-bar dx

OpenStudy (unklerhaukus):

\[\newcommand\p \newcommand \p \dd [1] { \,\mathrm d#1 } \p \pa [2] { \frac{\partial #1}{\partial #2} } \\ \begin{align*}f &=\int(3x^2y+6x)\dd x \\ &=x^3y+2x^2+g(y) \\ \\ \pa fy&=x^3+g'(y) =x^3+3y^2 \\ &\qquad g'(y) =3y^2 \\ &\qquad g(y) =y^3+c \\ \\ f(x,y) &=x^3y+2x^2+y^3+c \end{align*}\]

OpenStudy (unklerhaukus):

there is another mistakes somewhere

zepdrix (zepdrix):

\[\large \newcommand\p \newcommand \p \dd [1] { \,\mathrm d#1 } \p \pa [2] { \frac{\partial #1}{\partial #2} } \\ \begin{align*}f &=\int\limits(3x^2y+6x)\dd x \\ &=x^3y+\color{red}{2}x^2+g(y)\end{align*}\] Was this the mistake maybe? :o Looks like that 2 should be a 3.

OpenStudy (unklerhaukus):

ah yes. thankyou @zepdrix

zepdrix (zepdrix):

Hmm this looks like a fun problem. Grrr I gotta learn my maths :c

zepdrix (zepdrix):

You can has a medal also :3 You got through 98% of the problem without any trouble lol

OpenStudy (unklerhaukus):

Differential Equations are good fun.

OpenStudy (unklerhaukus):

now i am getting \[f(x,y)= x^3y+3x^2+y^3+c\\\,\\ x^3y+3x^2+y^3=k\]__________________ the solution in the back of the book say \[x^3y+3x^2y^3=c\] i guess they made a typo

OpenStudy (unklerhaukus):

if that's not fun i don't know what is